OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (jdoe0001):

\(\bf -3|2x + 6| = -12\implies |2x + 6| =\cfrac{-12}{-3}\implies|2x + 6| =4\\ \quad \\ |2x + 6| =4\implies \begin{cases} +(2x + 6) =4\\ \quad \\ \bf -(2x + 6) =4 \end{cases}\)

OpenStudy (jdoe0001):

so you would get 2 values for "x", one for each scenario

OpenStudy (anonymous):

-1 and -5 ?

OpenStudy (jdoe0001):

yeap